OpenStudy (anonymous):

What is the radius of a circle with a circumference of 23.55 meters? Use 3.14 for pi and show your work.

OpenStudy (anonymous):

please help me

OpenStudy (shane_b):

\[C=\pi d\]You know C and pi...and \[r=\frac{d}{2}\]so...what do you think you need to do?

OpenStudy (shane_b):

Or you could just blurt out the answer...which is *almost* correct.

OpenStudy (anonymous):

lol

OpenStudy (shane_b):

@yoyogators: Do you understand what to do at least? His answer only has a simple math error but the process is right.

OpenStudy (anonymous):

i dont trust him so can you give me step by step on how to get the correct answer?

OpenStudy (shane_b):

Ok...\[C=\pi d=\pi (2r)\]With me so far?

OpenStudy (anonymous):

Blurting out the answer with work :D

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

maybe you`ll need this: C = 2πr substitute the values 23.55=2(3.14)r 23.55=6.28(r) r=3.75m

OpenStudy (anonymous):

What is the radius of a circle with a circumference of 23.55 meters? Use 3.14 for pi and show your work.

OpenStudy (anonymous):

ok so that is the answer

OpenStudy (shane_b):

I think it's better to walk them through it and lead them to the answer...seeing the full answer right away is not always helpful.

OpenStudy (anonymous):

ok so they want to know what the radius is

OpenStudy (shane_b):

Right...and you know the circumference and pi. All this problem is is solving the formula for r.

OpenStudy (shane_b):

\[C=\pi(2r)\]So solving for r what do you get?

OpenStudy (anonymous):

so the r is half way in and if the circumference is 23.55

OpenStudy (shane_b):

yes...good so far.

OpenStudy (shane_b):

Forget the numbers for a moment and just solve the formula for r. Plugging in the values later is trivial.

OpenStudy (anonymous):

so you would divid 23.55 by 3.14 then divid by 2

OpenStudy (shane_b):

Correct. You end up with \[\frac{C}{2\pi}=r\]

OpenStudy (shane_b):

Now just plug in the values you were given.

OpenStudy (anonymous):

so if that is correct the raduis is 3.75

OpenStudy (shane_b):

Yes.

OpenStudy (anonymous):

so hold on is this correct

OpenStudy (anonymous):

You have to subtract the 1 first 4/5 and 5/6 Make the denominators equal 24/30, 25/30 Jake has the smaller fraction which means his time is lower.. Or divide top number by bottom and the smaller decimal wins. Divide 23.55 by pi, then divide by 2.

OpenStudy (anonymous):

opps pellet that is the wrong thing sorry

OpenStudy (anonymous):

S H I T not pellet

OpenStudy (shane_b):

Both of your assertions in the other question are correct. If the denominators are the same, the smaller one is the smaller amount. Or you can simply divide each fraction and the smaller decimal is the smaller value.

OpenStudy (anonymous):

ok so you would say to find what the radius you have to get 23.55 divid that by 3.14 then divid that by 2 and the answer will be 3.75

OpenStudy (anonymous):

ok so is what i wrote correct to the question

OpenStudy (shane_b):

Yes...or you could simplify more by saying: "Take the circumference and divide it by 2 times pi to get the radius"

OpenStudy (anonymous):

ok thanks man you are a life saver i will give you a metal!!!!!!!!

OpenStudy (shane_b):

No problem..good luck

OpenStudy (anonymous):

o i have one more sorry

OpenStudy (anonymous):

Which of the following is not a true statement? Pi is the ratio of a circle's circumference to its diameter. The center of a circle is the distance from one point on a circle to the opposite side. The radius of a circle has one endpoint at the center and the other endpoint on the circle. The circumference of a circle is twice the product of and its radius.

OpenStudy (shane_b):

Take a look at the formula: \[C=\pi2r\]Now solve it for pi:\[\pi=\frac{C}{2r}\] So which answer does that look like?

OpenStudy (shane_b):

Still here?

OpenStudy (anonymous):

ya sorry my mom called me

OpenStudy (anonymous):

ok let me look

OpenStudy (anonymous):

The center of a circle is the distance from one point on a circle to the opposite side.

OpenStudy (shane_b):

Ok, let's start with that. The question is stating that the center of a circle is the length of the line in the drawing: Is that true?

OpenStudy (anonymous):

yes

OpenStudy (shane_b):

No :) The length from one side of a circle to the opposite side is the diameter..so that statement is false.

OpenStudy (shane_b):

All of the other statements are true if you think about them.

OpenStudy (anonymous):

ok the wrong statement is The center of a circle is the distance from one point on a circle to the opposite side.

OpenStudy (shane_b):

Right

OpenStudy (anonymous):

ok thanks i will remember that now!

OpenStudy (shane_b):

k, good luck

OpenStudy (anonymous):